L Q AP MP

0 0 - -

1 8 8 8

2 22 11 14

3 60 20 38

4 96 24 36

5 132 26.4 26.4

6 132 22 0


The variable cost (VC) of the business consists of labor costs and raw material costs. The wage of labor is w = 3.000 monetary units and is fixed for each worker.

The cost of raw materials (c) is fixed and equal for each unit of product produced.

Find the fixed cost (FC) of the business if at the level of production of the 5th worker the marginal cost (MC) is 525 monetary units and the average total cost (ATC) is 700 monetary units.


1
Expert's answer
2021-11-29T15:12:43-0500

At the level of production of the 5th worker:

Q = 132 units, ATC = 700, so TC = 700*132 = 92,400.

"FC = TC - VC = 92,400 - (3 + c)*132 = 92,004 - 132c."